OpenStudy (anonymous):

Find the 6th term of the sequence: 8, 32, 128

OpenStudy (landshark5):

I got 8192

OpenStudy (anonymous):

so notice that each term after is multiplied by 4. 8 times 4 is 32. 32 times 4 is 128. 128 times 4 is 512. 512 times 4 is 2048. 2048 times 4 is 8192. The 6th term is 8192

OpenStudy (here_to_help15):

Do u still need help?

OpenStudy (anonymous):

I have another question

OpenStudy (here_to_help15):

ok

OpenStudy (anonymous):

find the 25th term of the sequence 22,26,30

OpenStudy (landshark5):

I got 122. I added 4 to each number.

OpenStudy (here_to_help15):

oh huh ok looks like @landshark5 gots it :P

OpenStudy (landshark5):

I am the human calculator. :D

OpenStudy (here_to_help15):

: P

OpenStudy (landshark5):

instead of adding i multiplied 4 by 25 and got 100 and i added the 100 to the 22, I got 122. Its faster than just adding.

OpenStudy (anonymous):

I got one more and this is my last question

OpenStudy (anonymous):

Toby was asked to find the sum of the first 10 terms of the arithmetic sequence but he plugged something into the formula wrong! Where did he make his mistake? What should he have done instead?

OpenStudy (anonymous):

7,4,1,-2...

OpenStudy (landshark5):

who's Toby? lol jk :3

OpenStudy (anonymous):

S10= 10/2 [2(7)+ (10-1)(3)] =5[14 + 9 (3)] = 5[14 + 27] =5[41] =205

OpenStudy (anonymous):

Can you help?

OpenStudy (landshark5):

ill try

OpenStudy (landshark5):

is the s a variable

OpenStudy (anonymous):

Please thank you so much

OpenStudy (anonymous):

S is the part of the formula

OpenStudy (landshark5):

i get it let me do the problem on paper

OpenStudy (anonymous):

ok

OpenStudy (landshark5):

I get 205 also. I can't find his mistake.

OpenStudy (landshark5):

wait one more try.

OpenStudy (anonymous):

Its like the order he did it or he plugged in the wrong thing

myininaya (myininaya):

hey is the common difference 3 or -3?

myininaya (myininaya):

look at your sequence of numbers it is decreasing

myininaya (myininaya):

the 10th term is in no way a bigger number than what you started with

OpenStudy (anonymous):

what do you mean

myininaya (myininaya):

7,4,1,-2... this is your sequence the numbers are getting smaller how could the 10th number be bigger than the first number? it can't

myininaya (myininaya):

do you see what I'm saying?

myininaya (myininaya):

basically this confirms you are using the wrong formula?

myininaya (myininaya):

or toby

myininaya (myininaya):

and that isn't a question